### v1

# Domain-specific Question Answering with Hybrid Search

Abstract
Domain-specific question answering is an evolving field that requires specialized solutions to address unique challenges. In this paper, we show that a hybrid approach—combining a fine-tuned dense retriever with keyword-based sparse search methods—significantly enhances performance. Our system leverages a linear combination of relevance signals, including cosine similarity from dense retrieval, BM25 scores, and URL host matching, each with tunable boost parameters. Experimental results indicate that this hybrid method outperforms our single-retriever system, achieving improved accuracy while maintaining robust contextual grounding. These findings suggest that integrating multiple retrieval methodologies with weighted scoring effectively addresses the complexities of domain-specific question answering in enterprise settings.
Introduction
With the increasing adoption of Large Language Models (LLMs) in enterprise settings, ensuring accurate and reliable question-answering systems remains a critical challenge. Building upon our previous work on domain-specific question answering about Adobe products (Sharma et al. 2024), which established a retrieval-aware framework with self-supervised training, we now present a production-ready, generalizable architecture alongside a comprehensive evaluation methodology. Our core contribution is a flexible, scalable framework built on Elasticsearch that can be adapted for any LLM-based question-answering system. This framework seamlessly integrates hybrid retrieval mechanisms, combining dense and sparse search with boost matching, while maintaining production-grade performance requirements. While we demonstrate its effectiveness using our organization’s domain-specific data, the architecture is designed to be domain-agnostic and can be readily deployed for other enterprise applications. We evaluate our system through a rigorous methodology that assesses performance across multiple dimensions: context relevance through normalized Discounted Cumulative Gain (nDCG), response groundedness, answer accuracy, and answer null rate. To ensure robust testing, we compile a diverse evaluation dataset including:
  1. 1.
    Human-annotated responses for common feature queries
  2. 2.
    A carefully curated “negative” dataset containing jailbreak attempts, NSFW content, and irrelevant queries to test system boundaries
  3. 3.
    LLM-based comparative analysis between system outputs and human-annotated ground truth
This comprehensive evaluation approach allows us to assess not only the system’s ability to provide accurate information but also its robustness against inappropriate queries and its capability to acknowledge when information is not available. Our hybrid search mechanism demonstrates significant improvements across these metrics compared to single-method approaches. The key contributions of this work include:
  • A production-ready, generalizable framework for LLM-based QA systems built on Elasticsearch
  • A flexible hybrid retrieval mechanism combining dense and sparse search methods
  • A comprehensive evaluation framework for assessing QA system performance
  • Empirical analysis demonstrating the effectiveness of our approach across various metrics
Through this work, we provide not only theoretical insights but also a practical, deployable solution for building reliable domain-specific question-answering systems that can be adapted to various enterprise needs.
Related Work
Our research builds upon recent advancements in retrieval-augmented language models, domain-specific question answering (QA), and hybrid retrieval methods. We review relevant literature across these key areas.
Retrieval-Augmented Language Models
Retrieval-Augmented Generation (RAG) models integrate information retrieval with language models to enhance performance on knowledge-intensive tasks such as question answering. These models employ a retriever to identify pertinent documents, conditioning the language model on the retrieved context during response generation. Lewis et al. (2020) introduced the RAG framework, demonstrating its effectiveness in open-domain QA by combining dense passage retrieval with sequence-to-sequence generation. Further, Lazaridou, Cancedda, and Baroni (2022) explored the application of RAG in an enterprise customer support setting, highlighting its potential for domain-specific QA.
Domain-Specific Question Answering
Developing QA systems tailored to specialized domains necessitates techniques that effectively capture domain-specific knowledge and terminology. Sharma et al. (2024) proposed a framework for compiling domain-specific QA datasets and introduced retrieval-aware fine-tuning of language models, which reduces hallucinations and enhances contextual grounding. Eppalapally et al. (2024) introduced a model for query rewriting to take the domain of interest into account. Additionally, hybrid retrieval methods that combine dense and sparse techniques have shown promise in domain-specific QA. For instance, Zhu et al. (2023) presented a hybrid text generation-based query expansion method, improving retrieval accuracy and QA system performance.
Hybrid Retrieval Methods
Integrating multiple retrieval signals can improve ranking performance compared to single-method approaches. Arivazhagan et al. (2023) explored hybrid hierarchical retrieval, combining sparse and dense methods in a two-stage document and passage retrieval setup, demonstrating improved in-domain and zero-shot generalization. Similarly, Li et al. (2021) proposed a dual reader-parser architecture that leverages both textual and tabular evidence, enhancing open-domain QA performance.
While prior work has made significant strides, developing reliable and robust domain-specific QA systems for enterprise settings remains an open challenge. Our work addresses this by proposing a flexible, generalizable framework that integrates hybrid retrieval with weighted relevance scoring. We extend prior hybrid approaches with a multi-phase scoring algorithm and introduce a comprehensive evaluation methodology to assess system performance across multiple dimensions.
Methodology
To optimize document retrieval accuracy, we employed a multi-phase scoring algorithm that evolved through iterations, starting from a baseline BM25 model, progressing to a chunking-based approach, and finally integrating host-based score adjustments.
Figure 1: Production ready RAG deployment
Scoring Algorithm
Given a user query, we calculate relevance scores for each document as follows:
[…]
Here, the final score combines the highest cosine similarity score from matched content chunks within each document, boosted BM25 scores, and a host-based weighting. Each document’s chunked cosine score ensures finer-grained relevance, while BM25 and host boosts adjust for term frequency relevance and content source authority, respectively. This combination provides a more nuanced ranking by integrating multiple retrieval signals.
Parameter Selection
The values for bm25_boost and host_boost were empirically determined by optimizing for the highest average similarity score between generated outputs and a predefined golden set. This iterative tuning process ensured that the scoring weights effectively balanced relevance and content source reliability. Table 2 and Table- 3 summarize how we select these parameters.
Document Ranking
After calculating scores for all documents, they were sorted in descending order, and the top 3 documents were returned as the final output, prioritizing the documents most likely to yield relevant information.
Experiments and Evaluation
Datasets
We evaluate our system using two carefully curated datasets: a golden dataset for measuring accuracy and performance, and a negative dataset for testing system robustness.
Golden Dataset
Our golden dataset comprises question-answer pairs from three primary documentation sources for our organization, with the following distribution:
  • creativecloud.adobe.com (Express Learn): 62 pairs
  • www.adobe.com: 74 pairs
  • helpx.adobe.com: 51 pairs
Each entry in the dataset contains:
  • A help query, initially LLM-generated and human-revised
  • Relevant document URLs (separated by delimiters)
  • Step-by-step answers, LLM-generated and human-revised
  • Source documentation URL
Negative Dataset
To evaluate system robustness, we compiled a negative dataset consisting of:
  • 12 jailbreak attempts
  • 6 NSFW queries
  • 12 irrelevant queries
All queries were collected from internet sources and revised by human annotators, with an expected ”content not found” response.
Evaluation Setup
Our evaluation framework uses the following common parameters across all experiments:
  • Generator Model: GPT-4o
  • Retrieved URLs Limit: 3
Figure 2: Evaluation Setup
Results and Analysis
Effect of Chunk Size on Retrieval Performance
To determine the optimal chunk size for document segmentation, we experimented with various chunk sizes and overlap values. Table 1 summarizes the NDCG scores obtained for different configurations.
Chunk size Chunk Overlap Context nDCG
1000 100 0.828
2000 500 0.802
5000 1000 0.795
Table 1: Effect of Approximate Chunk Size on Retrieval Performance
The results indicate that an approximate chunk size of 1000 characters with a split ensuring it ends on a sentence delimiter, with an overlap of 100 characters provides a favorable balance between retrieval granularity and computational efficiency.
Boost Parameter Tuning
We performed empirical tuning of the boost parameters to optimize the scoring algorithm for document retrieval.
BM25 Boost Tuning
Using 60% of the golden dataset as a validation set, we adjusted the BM25 content match boost. Table 2 shows the Top-3 NDCG scores for different boost values.
bm25-boost Context nDCG
0.1 0.863
0.3 0.868
0.6 0.840
1 0.831
Table 2: Choosing optimal parameter for BM25
The optimal BM25 content boost was found to be 0.3, achieving the highest NDCG score.
Host Boost Tuning
We also tuned the host boost parameter to enhance retrieval from authoritative sources. Table 3 presents the NDCG scores for varying host boost values.
host-boost Context nDCG
0.1 0.853
0.3 0.842
0.6 0.830
1 0.781
Table 3: Choosing optimal parameter for BM25
A host boost value of 0.1 provided the best trade-off between relevance and authority.
Discussion of Boost Tuning Results
The tuning of boost parameters significantly impacted retrieval performance. Higher boost values for BM25 content match improved NDCG up to a point, after which performance declined due to overemphasis on term frequency. Similarly, moderate host boosting enhanced the retrieval of authoritative documents without overshadowing relevance.
Retrieval Strategy
We evaluated five increasingly sophisticated retrieval strategies to assess their effectiveness in our domain-specific question-answering system. The strategies are:
  • Keyword Search (BM25): This baseline approach utilizes the BM25 algorithm for sparse keyword-based retrieval. It serves as a foundational benchmark for comparison.
  • OpenAI text-embedding-3-large: This strategy employs OpenAI’s pre-trained text-embedding-3-large model for dense retrieval. It leverages semantic embeddings to capture the meaning of queries and documents without domain-specific fine-tuning.
  • Fine-tuned Retriever: This approach uses a dense retriever that has been fine-tuned on our domain-specific dataset, as discussed briefly in the Methods section. The fine-tuning enhances the model’s ability to understand domain-specific terminology and nuances.
  • Fine-tuned Retriever + Keyword Search: This hybrid method combines our fine-tuned dense retriever with the BM25 keyword search. The integration aims to leverage both semantic understanding and exact keyword matching to improve retrieval performance.
  • Fine-tuned Retriever + Keyword Search + Host Boost Building upon the hybrid approach, this strategy incorporates URL host boosting. By assigning additional weight to documents from preferred URL hosts, we aim to enhance the relevance of the retrieved results in the enterprise context.
Strategy Context NDCG
Keyword Search (BM25) 0.640
text-embedding-3-large (openai) 0.760
Fine tuned retriever 0.828
Fine tuned retriever + Keyword Search 0.845
Fine tuned retriever + Keyword Search + Host boosting 0.847
Table 4: Detailed Performance Comparison of Retrieval Strategies
Guardrails around generated answer
In order to protect any jailbreak attempts, we have implemented a guardrail mechanism on top of answer generation. Before returning the generated answer to the user, we compare the similarity of the generated answer with the system prompt, excluding the user query. If this similarity is very high, we do not return any content to the user. We have added jailbreak attempt queries in the evaluation dataset as well, which proves that we have made the system robust with this additional check.
Generated Answer Analysis
To validate our hypothesis that a better retriever leads to better final responses, we conducted an analysis comparing the generated answers from the LLM to the golden answers using an LLM as an evaluator.
Strategy Answer Similarity mean Answer Similarity Std Groundedness mean Groundedness Std
Keyword Search (BM25) 0.717 0.27 0.919 0.27
text-embedding-3-large (openai) 0.748 0.23 0.979 0.13
Fine tuned retriever 0.755 0.22 0.974 0.15
Fine tuned retriever + Keyword Search 0.767 0.20 0.983 0.12
Fine tuned retriever + Keyword Search + Host boosting 0.780 0.19 0.983 0.12
Table 5: Performance Comparison of Final answer with different retrieval strategies
Methodology
Using GPT-4 as the evaluation model, we assessed the quality of the generated answers across different retrieval strategies. We focused on two key metrics:
Answer Similarity: Measures the degree of correspondence between the generated answer and the golden answer. Groundedness: Evaluates how well the generated answer is supported by the retrieved documents, ensuring factual correctness and context relevance. For each query in the golden dataset, we generated answers using the LLM with contexts retrieved by each strategy. We then prompted GPT-4 to rate the similarity and groundedness of each generated answer compared to the golden answer on a scale from 0 to 1.
Results
Table 5 presents the mean and std for answer similarity and groundedness across the different retrieval strategies.
The results indicate that strategies incorporating both the fine-tuned retriever and keyword search, especially with host boosting, achieved the highest scores in both answer similarity and groundedness.
Analysis
The incremental improvements in both answer similarity and groundedness scores align with our hypothesis that enhanced retrieval strategies lead to better final responses. Specifically:
  • Fine-tuned Retriever: By capturing domain-specific semantics, it provides more relevant context to the LLM, improving answer quality.
  • Hybrid Approach: Combining dense retrieval with keyword search leverages both semantic understanding and exact term matching, further enhancing the context.
  • Host Boosting: Prioritizing authoritative sources increases the reliability of the information provided, which is reflected in higher groundedness scores.
These findings demonstrate that optimizing retrieval not only improves retrieval metrics but also has a significant positive impact on the quality of the final answers generated by the LLM.
Discussion
Our comprehensive evaluation of the hybrid search framework reveals several key findings and implications for domain-specific question answering systems:
Performance Advantages of Hybrid Retrieval
The experimental results demonstrate that our hybrid approach, combining fine-tuned dense retrieval with sparse search and host boosting, consistently outperforms single-method approaches across multiple metrics. This superior performance can be attributed to several factors:
  • Complementary Retrieval Signals: The integration of dense and sparse retrieval methods allows the system to leverage both semantic understanding and exact keyword matching. While the fine-tuned dense retriever captures domain-specific context and terminology, the BM25-based keyword search ensures critical term matches aren’t overlooked.
  • Host Boosting Effect: The addition of host-based boosting further refined results by prioritizing authoritative sources, leading to a modest but consistent improvement in both retrieval accuracy (NDCG: 0.84701) and answer quality metrics (Answer Similarity: 0.78, Groundedness: 0.983).
Impact on Answer Generation
The improvement in retrieval quality shows a direct correlation with enhanced answer generation:
  • Answer Similarity: The progression from basic keyword search (0.717) to the full hybrid approach (0.78) demonstrates that better context retrieval leads to more accurate answers.
  • Groundedness: High groundedness scores across all methods indicate that the system maintains strong factual alignment, with the hybrid approach achieving the highest score (0.983).
System Robustness
Our evaluation framework’s inclusion of negative test cases (jailbreak attempts, NSFW queries, and irrelevant queries) demonstrates the system’s resilience to potential misuse. The implemented guardrail mechanism, which checks answer similarity against system prompts, provides an effective defense against jailbreak attempts while maintaining system usability.
Practical Implications
The findings have several important implications for enterprise deployments:
  • Scalability: The framework’s architecture, built on Elasticsearch, provides a production-ready solution that can handle enterprise-scale deployments while maintaining performance.
  • Adaptability: The tunable boost parameters (BM25_boost and host_boost) offer flexibility in adjusting the system for different domain-specific applications and content sources.
  • Cost-Effectiveness: The chunking optimization (optimal size: 1000 tokens with 100 token overlap) provides an effective balance between retrieval accuracy and computational efficiency.
These findings provide a foundation for future research in domain-specific question answering systems while offering practical insights for enterprise implementations.
Appendix A Model Fine-tuning Details
Retriever Model Configuration
We build a text encoder that computes a vector representation of a sentence. The model is trained to map user queries and corresponding documents in similar latent spaces so that it can be used as a semantic retriever (Reimers 2019).
Training Process
We use user behavior data to train the retriever. The behavior data consists of a ”user query” and a list of ”documents” that were shown to the user by the in-house system. Among the list of ”documents,” we have a signal that a user clicked on one of the provided documents. Since each document consists of a ”title” and ”body text,” we formulate the training loss as follows:
$L_{\text{total}}:=L_{1}\big{(}f_{\theta}(\text{q}),f_{\theta}(\text{title})\big{)}+L_{1}\big{(}f_{\theta}(\text{q}),f_{\theta}(\text{body text})\big{)},$
where $q$ is the user query, $f_{\theta}()$ is the sentence encoder parameterized by $\theta$, and $L_{1}$ represents the InfoNCE loss function that operates between images and text (Oord, Li, and Vinyals 2018).
Appendix B Limitations
While our hybrid search approach demonstrates strong performance for domain-specific question answering, several important limitations should be noted:
Context Awareness Limitations
The current system lacks crucial contextual awareness in several dimensions:
  1. 1.
    User Location Context: The system cannot differentiate between users on different pages or sections of the product (e.g., homepage versus editor), leading to potentially misaligned responses. This becomes particularly evident in answers that reference generic locations like “open homepage” without considering the user’s current position in the product journey.
  2. 2.
    Platform and Device Context: The system does not receive or account for information about:
    • Operating system platform (Windows, iOS, etc.)
    • Device type (mobile phone, desktop computer, tablet)
    • Screen size or resolution
    This limitation can result in instructions that may not be applicable to the user’s specific platform or device configuration.
Language Support
The current implementation is restricted to English-language queries and content. This represents a significant limitation for global enterprise deployment, particularly for multinational organizations requiring multilingual support. The system would need substantial modifications to:
  • Handle queries in multiple languages
  • Process multilingual documentation
  • Generate responses in the user’s preferred language
  • Account for cultural and regional differences in terminology
Multimodal Content Handling
A notable limitation exists in the system’s ability to process and leverage visual content:
  1. 1.
    Visual Context Loss: Many retrieved documents contain valuable screenshots, diagrams, or step-by-step visual guides that demonstrate product functionality. However, these visual elements are currently not:
    • Included in the context provided to the LLM
    • Used to enhance response generation
    • Incorporated into the scoring mechanism
  2. 2.
    Response Limitations: The system cannot:
    • Reference specific visual elements in responses
    • Generate or modify visual content
    • Provide visual feedback or confirmation
This limitation is particularly significant in software documentation, where visual guides often provide crucial information that text alone cannot effectively convey.
Future Work
Addressing these limitations suggests several promising directions for future research:
  1. 1.
    Enhanced Context Integration: Developing mechanisms to capture and utilize user context, including:
    • Real-time user location tracking within the product
    • Platform and device detection
    • Context-aware response generation
  2. 2.
    Multilingual Extension: Expanding the system to support multiple languages through:
    • Multilingual dense retriever training
    • Cross-lingual document retrieval
    • Language-specific response generation
  3. 3.
    Multimodal Enhancement: Incorporating visual content through:
    • Multimodal embedding models
    • Visual content understanding and extraction
    • Integration of visual elements in response generation
    • Screenshot and image-aware context processing
These improvements would significantly enhance the system’s utility and applicability across diverse enterprise environments.
Appendix C Evaluation Metrics Details
Normalized Discounted Cumulative Gain (nDCG)
nDCG is a standard metric in information retrieval that measures ranking quality. It is calculated as:
[…]
where DCG@k is defined as:
[…]
Here:
  • $rel_{i}$ is the relevance score of the result at position $i$
  • IDCG@k is the DCG@k value for the ideal ranking
  • The score ranges from 0 to 1, with 1 being perfect ranking
Interpretation of nDCG Scores
This metric is particularly suitable for our evaluation because:
  • It accounts for both relevance and position in the ranking
  • It is normalized, allowing comparison across queries
LLM-based Evaluation Metrics
In addition to nDCG, we employ GPT as an automated judge to evaluate two key aspects of the generated responses: groundedness and accuracy.
Groundedness Score
We evaluate how well the generated responses are grounded in the retrieved context using the following prompt structure (Truera 2023):
System: You are an impartial groundedness judge. You will be given a context and a response. Your task is to determine how grounded the response is in the given context. A response is considered grounded if it is supported by and does not contradict the given context. Rate the groundedness on a scale from 0 to 1 (0.0, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9, 1.0), where 0 is completely ungrounded and 1 is perfectly grounded. Context: \$context Response: \$response Groundedness Score:
Accuracy Score
We evaluate the accuracy of generated responses by comparing them with ground truth answers using the following prompt (Truera 2023):
System: You will be given one Model_answer and a Groundtruth_answer for a question about \$product. Your task is to rate the similarity of the two answers on one metric. Evaluation Criteria:
Similarity (0.0, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9, 1.0) - Similarity of overall text and similarity of each step or multiple steps also need to be considered.
Question: \$question Groundtruth_answer: \$ground_truth Model_answer: \$model_answer Evaluation Steps: 1. Read the question carefully 2. Do not modify the Groundtruth_answer and Model_answer 3. Read the Groundtruth_answer and identify the steps and order 4. Read the Model_answer and assess similarity to Groundtruth 5. Assign similarity score (0.0 to 1.0) 6. Reply with the Similarity score only
Both metrics output scores in the range [0,1], where:
  • Groundedness Score measures how well the response is supported by the retrieved context
  • Accuracy Score measures the semantic and structural similarity between the generated response and the ground truth
This combination of retrieval-based (nDCG) and generation-based (Groundedness, Accuracy) metrics provides a comprehensive evaluation framework for our system’s performance.

### v2

@@
 These findings suggest that integrating multiple retrieval methodologies with weighted scoring effectively addresses the complexities of domain-specific question answering in enterprise settings.
-Introduction
+1 Introduction
 With the increasing adoption of Large Language Models (LLMs) in enterprise settings, ensuring accurate and reliable question-answering systems remains a critical challenge.
@@
 Through this work, we provide not only theoretical insights but also a practical, deployable solution for building reliable domain-specific question-answering systems that can be adapted to various enterprise needs.
-Related Work
+2 Related Work
 Our research builds upon recent advancements in retrieval-augmented language models, domain-specific question answering (QA), and hybrid retrieval methods.
@@
 We extend prior hybrid approaches with a multi-phase scoring algorithm and introduce a comprehensive evaluation methodology to assess system performance across multiple dimensions.
-Methodology
+3 Methodology
 To optimize document retrieval accuracy, we employed a multi-phase scoring algorithm that evolved through iterations, starting from a baseline BM25 model, progressing to a chunking-based approach, and finally integrating host-based score adjustments.
@@
 After calculating scores for all documents, they were sorted in descending order, and the top 3 documents were returned as the final output, prioritizing the documents most likely to yield relevant information.
-Experiments and Evaluation
+4 Experiments and Evaluation
 Datasets
@@
 Effect of Chunk Size on Retrieval Performance
-To determine the optimal chunk size for document segmentation, we experimented with various chunk sizes and overlap values. Table 1 summarizes the NDCG scores obtained for different configurations.
+To determine the optimal chunk size for document segmentation, we experimented with various chunk sizes and overlap values.
+Table 1 summarizes the NDCG scores obtained for different configurations (see Appendix B for more details about the metric).
 Chunk size Chunk Overlap Context nDCG
@@
 It leverages semantic embeddings to capture the meaning of queries and documents without domain-specific fine-tuning.
-  • Fine-tuned Retriever: This approach uses a dense retriever that has been fine-tuned on our domain-specific dataset, as discussed briefly in the Methods section.
+  • Fine-tuned Retriever : As shown in Appendix A this approach uses a dense retriever that has been fine-tuned on our domain-specific dataset, as discussed briefly in the Methods section.
 The fine-tuning enhances the model’s ability to understand domain-specific terminology and nuances.
@@
 Using GPT-4 as the evaluation model, we assessed the quality of the generated answers across different retrieval strategies. We focused on two key metrics:
-Answer Similarity: Measures the degree of correspondence between the generated answer and the golden answer.
+Answer Similarity: Measures the degree of correspondence between the generated answer and the golden answer (see Appendix B).
 Groundedness: Evaluates how well the generated answer is supported by the retrieved documents, ensuring factual correctness and context relevance.
 For each query in the golden dataset, we generated answers using the LLM with contexts retrieved by each strategy.
-We then prompted GPT-4 to rate the similarity and groundedness of each generated answer compared to the golden answer on a scale from 0 to 1.
+We then prompted GPT-4 to rate the similarity and groundedness of each generated answer compared to the golden answer on a scale from 0 to 1 (see Appendix B).
 Results
@@
   • Host Boosting: Prioritizing authoritative sources increases the reliability of the information provided, which is reflected in higher groundedness scores.
+Human-GPT Score Correlation Analysis
+To validate our LLM-based evaluation approach, we performed a focused correlation analysis using a subset of 14 queries that appeared in both golden dataset and a larger human evaluation dataset (containing approximately 200 queries).
+For these overlapping queries, we compared GPT-4’s answer accuracy scores with binary human feedback (1 for correct, 0 for incorrect).
+While this represents a limited sample, it provides valuable initial insights into the alignment between automated and human evaluation methods.
+Analysis of these scores revealed:
+  • Strong positive correlation between GPT-4 and human evaluations (Pearson correlation coefficient = $0.89$)
+  • GPT-4’s scores showed good discrimination, with higher scores ($\mu=0.82$) for queries rated correct by humans
+  • Only one case of significant disagreement occurred (for the Generative Fill query)
+While this analysis is based on a small overlapping set, the strong correlation suggests promise in using LLM-based evaluation as a complement to human assessment.
+Future work should focus on expanding this comparison using a larger set of consistently evaluated queries to more comprehensively validate the automated evaluation approach.
 These findings demonstrate that optimizing retrieval not only improves retrieval metrics but also has a significant positive impact on the quality of the final answers generated by the LLM.
@@
 Our evaluation framework’s inclusion of negative test cases (jailbreak attempts, NSFW queries, and irrelevant queries) demonstrates the system’s resilience to potential misuse.
+Detailed analysis of system performance against potentially problematic queries revealed:
+Query Type Performance Details
+Jailbreak Attempts 91.7% 11/12 null responses
+NSFW Content 100% 6/6 null responses
+Irrelevant Queries 100% 12/12 null responses
+Table 6: System Performance on Negative Queries
+The system demonstrated particularly strong performance in handling NSFW content and irrelevant queries, with perfect accuracy in both categories.
+For jailbreak attempts, only one query out of twelve received any response, demonstrating robust resilience against prompt injection attempts.
 The implemented guardrail mechanism, which checks answer similarity against system prompts, provides an effective defense against jailbreak attempts while maintaining system usability.
@@
 These findings provide a foundation for future research in domain-specific question answering systems while offering practical insights for enterprise implementations.
+Future Work
+Comprehensive Human Evaluation Framework
+Building upon our initial correlation analysis between LLM and human evaluations, future work should prioritize:
+  • Development of a standardized evaluation protocol with consistent query sets for both human and automated evaluation, clear criteria for human evaluators and structured feedback collection mechanisms
+  • Large-scale human evaluation studies including systematic assessment of system performance across diverse query types, evaluation by both domain experts and end-users, collection of qualitative feedback on answer quality and usefulness and analysis of user satisfaction across different use cases
+Enhanced Context Integration
+: Developing mechanisms to capture and utilize user context, including:
+  • Real-time user location tracking within the product
+  • Platform and device detection
+  • Context-aware response generation
+Multilingual Extension
+: Expanding the system to support multiple languages through:
+  • Multilingual dense retriever training
+  • Cross-lingual document retrieval
+  • Language-specific response generation
+Multimodal Enhancement
+: Incorporating visual content through:
+  • Multimodal embedding models
+  • Visual content understanding and extraction
+  • Integration of visual elements in response generation
+  • Screenshot and image-aware context processing
+These improvements would significantly enhance the system’s utility and applicability across diverse enterprise environments.
 Appendix A Model Fine-tuning Details
@@
 where $q$ is the user query, $f_{\theta}()$ is the sentence encoder parameterized by $\theta$, and $L_{1}$ represents the InfoNCE loss function that operates between images and text (Oord, Li, and Vinyals 2018).
-Appendix B Limitations
-While our hybrid search approach demonstrates strong performance for domain-specific question answering, several important limitations should be noted:
-Context Awareness Limitations
-The current system lacks crucial contextual awareness in several dimensions:
-  1. 1.
-    User Location Context: The system cannot differentiate between users on different pages or sections of the product (e.g., homepage versus editor), leading to potentially misaligned responses.
-This becomes particularly evident in answers that reference generic locations like “open homepage” without considering the user’s current position in the product journey.
-  2. 2.
-    Platform and Device Context: The system does not receive or account for information about:
-    • Operating system platform (Windows, iOS, etc.)
-    • Device type (mobile phone, desktop computer, tablet)
-    • Screen size or resolution
-    This limitation can result in instructions that may not be applicable to the user’s specific platform or device configuration.
-Language Support
-The current implementation is restricted to English-language queries and content.
-This represents a significant limitation for global enterprise deployment, particularly for multinational organizations requiring multilingual support.
-The system would need substantial modifications to:
-  • Handle queries in multiple languages
-  • Process multilingual documentation
-  • Generate responses in the user’s preferred language
-  • Account for cultural and regional differences in terminology
-Multimodal Content Handling
-A notable limitation exists in the system’s ability to process and leverage visual content:
-  1. 1.
-    Visual Context Loss: Many retrieved documents contain valuable screenshots, diagrams, or step-by-step visual guides that demonstrate product functionality.
-However, these visual elements are currently not:
-    • Included in the context provided to the LLM
-    • Used to enhance response generation
-    • Incorporated into the scoring mechanism
-  2. 2.
-    Response Limitations: The system cannot:
-    • Reference specific visual elements in responses
-    • Generate or modify visual content
-    • Provide visual feedback or confirmation
-This limitation is particularly significant in software documentation, where visual guides often provide crucial information that text alone cannot effectively convey.
-Future Work
-Addressing these limitations suggests several promising directions for future research:
-  1. 1.
-    Enhanced Context Integration: Developing mechanisms to capture and utilize user context, including:
-    • Real-time user location tracking within the product
-    • Platform and device detection
-    • Context-aware response generation
-  2. 2.
-    Multilingual Extension: Expanding the system to support multiple languages through:
-    • Multilingual dense retriever training
-    • Cross-lingual document retrieval
-    • Language-specific response generation
-  3. 3.
-    Multimodal Enhancement: Incorporating visual content through:
-    • Multimodal embedding models
-    • Visual content understanding and extraction
-    • Integration of visual elements in response generation
-    • Screenshot and image-aware context processing
-These improvements would significantly enhance the system’s utility and applicability across diverse enterprise environments.
-Appendix C Evaluation Metrics Details
+Appendix B Evaluation Metrics Details
 Normalized Discounted Cumulative Gain (nDCG)